OpenStudy (anonymous):

rewrite the equation using function notation. x+1y=3 MEDALLLLLSLLSLSSLSLLSLSLSSS

OpenStudy (anonymous):

x + 1y = 3 x + y = 3 y = 3 -x [ f(x) = y] f(x) = 3 - x THAT'S IT !!!

OpenStudy (anonymous):

:D